solve 4x^(2) + 2 = 102 find x
Answer by stanbon(75887)

solve 4x^(2) + 2 = 102
4x^2 = 100
x^2 = 25
x^2-25 = 0
(x-5)(x+5) = 0
x = 5 or x = -5
=====================
Cheers,
Stan H.
==================